OpenStudy (anonymous):

PLEASE HELP ! I Will Give Metals ! What polynomial has roots of -6, -4, and 1?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@radar

OpenStudy (phi):

roots of -6, -4, and 1 means x= -6, x=-4 and x=1 or x+6 = 0, x+4=0 and x-1=0 multiply all together: (x+6)(x+4)(x-1)=0 that is your polynomial. multiply it out to put it into "standard form"

OpenStudy (anonymous):

so i'd do distributive property right now?

OpenStudy (anonymous):

I got x3 + x2 - 26x + 24, am i right?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

if you multiply just the numbers part you get 6*4*-1= -24 I would expect a -24 in the answer.

OpenStudy (anonymous):

hmm, so x3 - x2 - 26x - 24 ?

OpenStudy (phi):

what do you get for (x+6)(x+4) ?

OpenStudy (anonymous):

x2 + 4x + 6x + 24

OpenStudy (anonymous):

@phi

OpenStudy (phi):

now combine 4x + 6 x 4 x's plus 6 x's is how many x's ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

10x

OpenStudy (anonymous):

so x2 + 10x + 24 ?

OpenStudy (phi):

yes. now multiply by (x-1) \[ (x^2 +10x+24)(x-1) \]

OpenStudy (anonymous):

i got x3 - x2 + 10x - 10 + 24x + 24

OpenStudy (anonymous):

and then after combining like terms

OpenStudy (phi):

you should re-do it. If we distribute, we can write it like this \[ (x^2 +10x+24)(x-1) \\ x^2(x-1) +10x(x-1) + 24(x-1) \]

OpenStudy (anonymous):

okay, i thought i was doing something wrong

OpenStudy (phi):

you get \[ x^3 - x^2 + 10x\cdot x + 10x \cdot -1 + 24x + 24 \cdot -1 \] or \[ x^3 - x^2 + 10x^2 -10x +24x -24\] now combine like terms

OpenStudy (anonymous):

thats exactly what i got ! ^_^

OpenStudy (anonymous):

i got x3 + 9x2 + 14x - 24

OpenStudy (phi):

yes, that looks good

OpenStudy (anonymous):

Awesome, thanks ! could you possibly help me with one more?

OpenStudy (phi):

you must have posted a typo up above, because you had +24 the first time.

OpenStudy (anonymous):

oh i hadn't noticed that. I guess i did.

OpenStudy (anonymous):

could you help with one more?

OpenStudy (phi):

please make it a new post.

OpenStudy (anonymous):

okay